write an equation in slope intercept form of a linear function f whose graph passes through (-5,6) and is perpendicular to the line that has an x-intercept of 2 and a y-intercept of 9
Answer by lwsshak3(11628)

write an equation in slope intercept form of a linear function f whose graph passes through (-5,6) and is perpendicular to the line that has an x-intercept of 2 and a y-intercept of 9
**
Standard form of equation for a straight line: y=mx+b, m=slope, b=y-intercept
x-intercept is point (2,0)
y-intercept is point (0,9)
These two points are on a straight line,
Using two points above to find the slope m
m=∆y/∆x=(9-0)/(0-2)=-9/2
..
slope of linear function f, perpendicular to line=2/9 (negative reciprocal)
Equation: y=(2/9)x+b
solving for b using given point (-5,6) thru which f passes
6=(2/9)*-5+b
6=-10/9+b
b=6+10/9=54/9+10/9=64/9
Equation of f: y=2x/9+64/9